OpenStudy (anonymous):

Find the solutions in the interval [0, 2π). 3 tan^3 θ − 3 tan^2 θ − tan θ + 1 = 0

OpenStudy (anonymous):

one is tan thita =1 (by hit and trial method)

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

just put tan thita = 1 and u will realise that it is indeed a solution to the eqn

OpenStudy (anonymous):

then to find the other 2 solutions just divide the equation with tan thita -1 and get a quadratic in tan thita, which will give the other two roots

myininaya (myininaya):

\[3u^3-3u^2-u+1=0\] \[3u^2(u-1)-1(u-1)=0\] \[(u-1)(3u^2-1)=0\] \[=> u=1 \text{ or } u=\pm \sqrt{\frac{1}{3}}\]

myininaya (myininaya):

\[\tan(\theta)=1 \text{ or } \tan(\theta)=\pm \frac{1}{\sqrt{3}}\]